# Rethinking Pan-sharpening: Principled Design, Unified Training, and a Universal Loss Surpass Brute-Force Scaling

### v2

@@ Abstract
The field of pan-sharpening has recently seen a trend towards increasingly large and complex models, often trained on single, specific satellite datasets. This approach, however, leads to high computational overhead and poor generalization on full resolution data, a paradigm we challenge in this paper. In response to this issue, we propose PanTiny, a lightweight, single-step pan-sharpening framework designed for both efficiency and robust performance. More critically, we introduce multiple-in-one training paradigm, where a single, compact model is trained simultaneously on three distinct satellite datasets (WV2, WV3, and GF2) with different resolution and spectral information. Our experiments show that this unified training strategy not only simplifies deployment but also significantly boosts generalization on full-resolution data. Further, we introduce a universally powerful composite loss function that elevates the performance of almost all of models for pan-sharpening, pushing state-of-the-art metrics into a new era. Our PanTiny model, benefiting from these innovations, achieves a superior performance-to-efficiency balance, outperforming most larger, specialized models. Through extensive ablation studies, we validate that principled engineering in model design, training paradigms, and loss functions can surpass brute-force scaling. Our work advocates for a community-wide shift towards creating efficient, generalizable, and data-conscious models for pan-sharpening. The code is available at https://github.com/Zirconium233/PanTiny
@@ 1 Introduction
Pan-sharpening, a fundamental image fusion task in remote sensing, aims to merge a high-resolution panchromatic (PAN) image with a lower-resolution multispectral (LRMS) image to generate a single high-resolution multispectral (HRMS) image. This fused image is crucial for numerous downstream applications, including land-cover classification, environmental monitoring, and urban planning (Masi et al. 2016; Yang et al. 2017). Early approaches were dominated by traditional methods such as Component Substitution (CS) (Carper et al. 1990; Chavez and Kwarteng 1989) and Multi-Resolution Analysis (MRA) (King and Wang 2001; Liu 2000), which, while efficient, often introduced spectral and spatial distortions due to their handcrafted prior. The advent of deep learning, particularly with models like PNN (Masi et al. 2016) and PanNet (Yang et al. 2017), revolutionized the field by learning complex mappings directly from data and significantly improving fusion quality.
However, the recent pursuit of higher performance has led to a problematic trend. State-of-the-art (SOTA) methods increasingly rely on massive, complex architectures. For instance, CFDCNet (Li et al. 2025) achieves high metrics but at the cost of an astounding 55G FLOPs under $\times 128$ 128 resolution images. Other models like Pan-Mamba (He et al. 2025) perform well but show limited generalization capabilities. Methods such as PanFlow (Yang et al. 2023) and diffusion-based models (Zhong et al. 2025) are often multi-step, complicating the inference process. Conversely, lightweight solutions like Pan-LUT (Cai et al. 2025), while fast, exhibit a noticeable performance gap.
These specific shortcomings point to deeper, community-wide issues rooted in the pursuit of benchmark leadership over practical utility. A prevailing belief in “brute-force scaling” has led to models that are excessively large for the domain’s relatively small datasets (Deng et al. 2022). This issue is worsened by the community standard of training separate models for each satellite—a “one-dataset, one-model” philosophy that hinders both deployment efficiency and generalization. This reveals a critical issue: the majority of training for single-dataset models may contribute little to true, transferable generalization. This is evidenced by a finding in our appendix: a model fully converged on a source dataset shows nearly the same cross-domain performance as a simple baseline trained on that same source dataset for just one epoch.
In this work, we propose a comprehensive solution that challenges these norms. We first introduce PanTiny, a lightweight model that, as illustrated in Figure 1, strikes a good balance between performance and efficiency. Our extensive architectural ablations show that while scaling our model can further boost performance, it comes with diminishing returns, justifying our focus on efficiency. Second, we introduce a new “all-in-one” training paradigm in pan-sharpening domain, through extensive experiments, we demonstrate that this approach not only simplifies the deployment pipeline but also significantly improves generalization on full-resolution data. We believe our findings support a shift towards unified models. This direction is inspired by the foundation model trend in other vision fields, suggesting a promising path for creating more generalizable pan-sharpening solutions. Finally, we design a powerful and universal composite loss function that significantly elevates the performance of all tested models, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era.
@@
  • We propose and validate a new “all-in-one” training paradigm in pan-sharpening domain, demonstrating its ability to improve generalization on full-resolution data across various models.
  • We propose PanTiny, a lightweight yet powerful model that strikes a good balance between performance and efficiency, whose design is methodically guided by key insights distilled from our extensive architectural ablations.
  • We introduce a universally effective composite loss function that consistently boosts the performance of diverse models, setting a new benchmark for the pan-sharpening field.
@@ 2 Related Work
2.1 Traditional Pan-sharpening
Traditional pan-sharpening methods are generally categorized into Component Substitution (CS), Multi-Resolution Analysis (MRA), and hybrid approaches. CS-based methods, such as Intensity-Hue-Saturation (IHS) (Carper et al. 1990) and Principal Component Analysis (PCA) (Chavez and Kwarteng 1989), project the MS image into a different space, replace one component with the PAN image, and then perform an inverse transform. These methods excel at enhancing spatial details but often suffer from significant spectral distortion. MRA-based methods, like those using wavelet transforms (King and Wang 2001) or smoothing filters (e.g., SFIM (Liu 2000)), decompose the images into different frequency bands and inject the high-frequency details from the PAN image into the MS image. MRA methods generally preserve spectral information better but can introduce spatial artifacts.
2.2 Deep Learning-based Pan-sharpening
The success of deep learning in computer vision spurred its application in pan-sharpening. PNN (Masi et al. 2016) was a pioneering work that used a simple three-layer CNN to learn the mapping from up-sampled MS and PAN images to the high-resolution MS output. PanNet (Yang et al. 2017) improved upon this by working in the high-frequency domain and introducing a spectral loss to better preserve color information. Subsequent works explored more complex CNN architectures, such as MSDCNN (Yuan et al. 2018), which used multi-scale features to improve fusion quality. These methods consistently outperformed traditional techniques, setting a new standard for the field.
2.3 Recent Advances and SOTA Models
The current landscape of pan-sharpening is dominated by advanced deep learning architectures. Inspired by successes in other vision tasks, researchers have incorporated Transformers, State-Space Models (SSMs) like Mamba (He et al. 2025), and flow-based models (Yang et al. 2023). For instance, Pan-Mamba leverages the efficiency of SSMs to achieve impressive results. The very recent CFDCNet (Li et al. 2025) has pushed performance metrics to new heights, but at the cost of an enormous computational load (55G FLOPs). Other approaches, such as PSCINN (Wang et al. 2024), utilize invertible neural networks to model the fusion process. While powerful, these models often come with a substantial increase in parameters and complexity. Furthermore, a common thread among these SOTA methods is their training protocol: they are almost exclusively trained and tested on a single dataset. Some works explore generalization by training on one dataset and testing on others (Zhang et al. 2024), but the performance drop is often significant. The concept of an “all-in-one” model, trained jointly on multiple datasets for a single task, remains largely unexplored in pan-sharpening, representing a key opportunity that our work addresses.
@@ 3 Methodology
Our proposed method, PanTiny, is built on the principles of efficiency, simplicity, and empirical validation. We deliberately avoid overly complex operators and instead focus on a clean, effective architecture where each component’s inclusion is justified by extensive experiments. The overall architecture, shown in Figure 2, features a single-encoder design, a Transformer-based body for feature processing, and a simple convolutional refinement head.
@@ 3.1 Overall Architecture
Single Encoder
Unlike many methods that use separate encoders for PAN and MS inputs, we adopt an efficient single-encoder architecture. The upsampled MS image is first passed through a lightweight convolutional block to extract initial features. The PAN image is then integrated directly in the feature space via our fusion module. This design is highly parameter-efficient and, as our experiments show, forms the basis of a powerful and generalizable model.
Feature Fusion and Processing
Our investigation into fusion mechanisms revealed a surprising insight: in the multi-dataset training context, simplicity triumphs over complexity. We found that complex fusion strategies like cross-attention or the multi-layer “deepfusion” block from actually degraded performance compared to a simple baseline. We attribute this to overfitting. Complex fusion modules tend to memorize dataset-specific artifacts. This “specialized knowledge” fails to generalize when the model is required to perform across multiple datasets, whereas a simpler module is forced to learn more robust, common features. The core of our network is a series of standard Transformer blocks, which effectively model long-range dependencies and perform deep feature interaction. After the Transformer body, we use a simple fusion block composed of two consecutive 3x3 convolutional layers (“Enhanced Conv”). This design choice, validated in Table 7, proved to be the most effective and robust across all datasets.
Refinement Module
For the final reconstruction, we employ a single convolutional layer to map the fused features back to the desired high-resolution MS image. Our experiments (Table 8) confirmed that more elaborate refinement modules, such as those incorporating residual blocks or attention, offered no significant benefit and unnecessarily increased model size.
@@ 3.2 Transformer Block
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
Traditional Methods
Brovey (Gillespie, Kahle, and Walker 1987) - - 35.86 0.9216 0.0403 22.50 0.5466 0.1159 37.79 0.9026 0.0218
IHS (Carper et al. 1990) - - 35.29 0.9027 0.0461 22.55 0.5354 0.1266 38.17 0.9100 0.0243
SFIM (Liu 2000) - - 34.12 0.8975 0.0439 21.82 0.5457 0.1208 36.90 0.8882 0.0318
GS (Laben and Brower 2000) - - 35.63 0.9176 0.0423 22.56 0.5470 0.1217 37.22 0.9034 0.0309
Deep Learning Methods (All-in-One Training)
PNN (Masi et al. 2016) 68.9 2.26 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
PanNet (Yang et al. 2017) 80.3 2.63 38.98 0.9468 0.0301 29.12 0.8927 0.0935 43.26 0.9668 0.0176
MSDCNN (Yuan et al. 2018) 239.0 7.83 40.31 0.9580 0.0267 29.63 0.9033 0.0833 43.21 0.9671 0.0176
PanFlow (Yang et al. 2023) 87.3 2.86 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
PSCINN (Wang et al. 2024) 3321.5 108.84 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Pan-Mamba (He et al. 2025) 488.8 16.02 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
CFDCNet (Li et al. 2025) 1700.8 55.73 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
Figure 3: Quality comparison across SOTA methods on WV3 dataset. Refer to appendixs for more results.
@@ 3.3 Loss Function
  • L1 Loss: We use the Charbonnier loss (Charbonnier et al. 1994), a differentiable variant of the L1 norm that is less sensitive to outliers. For a batch of $B$ images with $N$ pixels each, it is defined as:
@@
  • SSIM Loss: To preserve perceptual quality and high-frequency structural details, we incorporate the Structural Similarity (SSIM) loss (Wang et al. 2004). The SSIM index between two image patches $o$ and $g$ is:
@@
  • Focal Loss for Regression: Inspired by focal loss in classification, we adapt it for regression to prioritize “hard” pixels that are more difficult to reconstruct. Let $d_{i}=|O_{i}-G_{i}|$ be the absolute error for a given pixel $i$. Our regression-style focal loss is formulated as:
@@
    where $r_{1}$ is a focusing parameter. This formulation up-weights pixels with larger errors, compelling the model to focus on challenging details.
@@
Through extensive experiments (see Table 2), we determined the optimal weights to be $\lambda_{1}=1.5$, $\lambda_{2}=4.0$, and $\lambda_{3}=1.5$, which consistently delivered the best performance across all datasets and models.

### v3

# Rethinking Pan-sharpening: A New Training Process for Full-Resolution Generalization

@@ Abstract
The field of pan-sharpening has recently seen a trend towards increasingly large and complex models, often trained on single, specific satellite datasets. This “one-dataset, one-model” approach leads to high computational overhead and impractical deployment. More critically, it overlooks a core challenge: poor generalization from reduced-resolution (RR) training to real-world full-resolution (FR) data. In response to this issue, we challenge this paradigm. We introduce a “multiple-in-one” training strategy, where a single, compact model is trained simultaneously on three distinct satellite datasets (WV2, WV3, and GF2). Our experiments show the primary benefit of this unified strategy is a significant and universal boost in FR generalization (QNR) across all tested models, directly addressing this overlooked problem. This paradigm also inherently solves the ’one-model-per-dataset’ challenge, and we support it with a highly reproducible, dependency-free codebase for true usability. Finally, we propose PanTiny, a lightweight framework designed specifically for this new, robust paradigm. We demonstrate it achieves a superior performance-to-efficiency balance, proving that principled, simple and robust design is more effective than brute-force scaling in this practical setting. Our work advocates for a community-wide shift towards creating efficient, deployable, and truly generalizable models for pan-sharpening. The code is open-sourced at https://github.com/Zirconium233/PanTiny .
@@ 1 Introduction
Pan-sharpening, a fundamental image fusion task in remote sensing, aims to merge a high-resolution panchromatic (PAN) image with a lower-resolution multispectral (LRMS) image to generate a single high-resolution multispectral (HRMS) image [13, 18]. This fused image is crucial for numerous downstream applications [13, 18]. Early approaches were dominated by traditional methods [3, 5, 9, 12]. The advent of deep learning, particularly with models like PNN [13] and PanNet [18], revolutionized the field by learning complex mappings directly from data and significantly improving fusion quality.
However, the recent pursuit of higher performance has led to a research landscape heavily focused on model architecture innovation. This trend generally follows three main directions. First, some works attempt to solve cross-dataset generalization, developing complex features to bridge the significant domain gaps between satellites [21]. Second, a dominant trend is the adaptation of SOTA architectures from other vision fields to pan-sharpening, including porting State-Space Models (SSMs) like Pan-Mamba [8] and flow-based models [17]. Third, some models achieve high metrics through brute-force scaling, resulting in massive, complex architectures. For instance, CFDCNet [11] achieves high metrics but at the cost of an astounding 55G FLOPs.
While these approaches have pushed benchmarks, they share common drawbacks: they are almost all evaluated on reduced-resolution (RR) benchmarks, and they adhere to a “one-dataset, one-model” philosophy. We argue this focus on architectural novelty overlooks two more fundamental, practical challenges. First, the “one-dataset, one-model” paradigm is impractical. It requires training and maintaining separate, specialized models for each satellite sensor, complicating real-world deployment. Second, and most critically, the true, practical challenge is full-resolution (FR) generalization—the ability for a model trained on RR data to perform well on real-world, full-resolution images. This crucial aspect is largely unaddressed. Furthermore, as our supplementary material shows, the domain gaps between satellites are so large that cross-dataset generalization (the goal of [21]) is often futile, suggesting that joint training is a more pragmatic path forward.
In this work, we propose a comprehensive solution that challenges these norms by shifting the focus from model architecture to the training process and practical efficiency. First, we introduce a new “all-in-one” training paradigm, representing a novel process-level innovation. We are the first to demonstrate that a single model can be jointly trained on multiple heterogeneous datasets (WV2, WV3, GF2). The primary benefit is not maximizing RR metrics, but as we will show, dramatically and universally boosting FR generalization (Table 2) across all tested models.
Second, this paradigm inherently solves the deployment and usability problem. As illustrated in Figure 1, a single, compact model can now service three distinct datasets, replacing the cumbersome ‘one-model-per-dataset’ approach.
Finally, we demonstrate that principled, lightweight engineering is superior to brute-force scaling. We propose PanTiny, a lightweight model designed specifically for this paradigm, and a universally powerful composite loss function. Our PanTiny model, benefiting from these innovations, achieves a superior performance-to-efficiency balance, outperforming most larger, specialized models [11]. Critically, unlike models such as Pan-Mamba [8] or CFDCNet [11], PanTiny uses only standard, universally available operators, making it dependency-free, easy to reproduce, and highly practical for deployment.
@@
  • We propose and validate a new “all-in-one” training paradigm, a process-level innovation that shifts the field’s focus from model-centric design and demonstrably boosts generalization on full-resolution data.
  • We address the ‘one-model-per-dataset’ problem by training a single, unified model for multiple datasets, significantly improving deployment efficiency and usability, as illustrated in Figure 1.
  • We propose PanTiny, a lightweight yet powerful model, and a universal composite loss, which together prove that principled, efficient design—which is also dependency-free and easy to reproduce—can surpass brute-force scaling within this new paradigm.
@@ 2 Related Work
2.1 Traditional and Early Deep Learning Methods
Traditional pan-sharpening methods include Component Substitution (CS) [3, 5], which enhances spatial detail but risks spectral distortion, and Multi-Resolution Analysis (MRA) [9, 12], which better preserves spectral information but can add spatial artifacts. Deep learning (DL) methods later surpassed these, with pioneering works like PNN [13] using a simple CNN. PanNet [18] improved upon this by working in the high-frequency domain, and subsequent models like MSDCNN [19] explored multi-scale features to enhance fusion quality.
2.2 SOTA Architectural Innovation
Recent work is dominated by model-centric innovation, often adapting SOTA architectures from other vision fields. This includes porting State-Space Models (e.g., Pan-Mamba [8]), flow-based models (e.g., PanFlow [17]), and invertible neural networks (e.g., PSCINN [15]). Another trend is brute-force scaling, exemplified by CFDCNet [11], which achieves high metrics but at an enormous computational cost (55G FLOPs).
2.3 Data Generalization Paradigms
Research on data generalization is also prevalent. Most methods follow a “one-dataset, one-model” philosophy. While some attempt cross-dataset generalization (e.g., train on A, test on B) by learning domain-irrelevant features [21], this approach often fails due to large domain gaps between sensors. The concept of an “all-in-one” model, trained jointly on multiple datasets, remains largely unexplored. Our work addresses this gap, proposing a unified paradigm to solve the deployment problem and, critically, to improve full-resolution generalization—a practical challenge largely unaddressed by the community.
@@ 3 Methodology
Our proposed method, PanTiny, is built on the principles of efficiency, simplicity, and empirical validation, designed specifically to be robust and effective within our “all-in-one” training paradigm. As our new focus is on multi-domain training, we deliberately avoid overly complex operators that, as we will show (Table 7), are prone to overfitting. We instead focus on a clean, effective architecture where each component’s inclusion is justified by extensive experiments. The overall architecture, shown in Figure 2, features a single-encoder design, a Transformer-based body for feature processing, and a simple convolutional refinement head.
@@ 3.1 Overall Architecture
3.1.1 Patch Embedding and Encoding
Unlike many methods that use separate encoders, we adopt an efficient single-encoder architecture. The upsampled MS image (bms) is first passed through an OverlapPatchEmbed layer (a 3x3 convolution) to extract initial features. These features are then processed by a shallow set of Transformer blocks (self.encoder in our implementation).
3.1.2 Feature Fusion Module
Following the initial encoding, the PAN image is integrated directly in the feature space via our fusion module. Our investigation into fusion mechanisms revealed a surprising insight: in the multi-dataset “all-in-one” context, simplicity triumphs over complexity. We found that complex, specialized fusion strategies like the multi-layer “deepfusion” block [8] (Table 7) degraded performance, suggesting overfitting to dataset-specific artifacts. This ”specialized knowledge” fails when the model must perform across multiple domains. Therefore, we adopt a simple but highly effective Enhanced Conv” block (two consecutive 3x3 convolutional layers), which our ablations (Table 7) proved to be the most robust.
3.1.3 Deep Feature Processing and Refinement
The fused features are then processed by the main body of the network (self.refinement in code), which consists of a deeper stack of standard Transformer blocks for effective long-range dependency modeling. For the final reconstruction, we again prioritize simplicity. Our experiments (Table 8) confirmed that a single convolutional layer was optimal, while more elaborate refinement modules offered no significant benefit and unnecessarily increased model size.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
Traditional Methods
Brovey [7] - - 35.86 0.9216 0.0403 22.50 0.5466 0.1159 37.79 0.9026 0.0218
IHS [3] - - 35.29 0.9027 0.0461 22.55 0.5354 0.1266 38.17 0.9100 0.0243
SFIM [12] - - 34.12 0.8975 0.0439 21.82 0.5457 0.1208 36.90 0.8882 0.0318
GS [10] - - 35.63 0.9176 0.0423 22.56 0.5470 0.1217 37.22 0.9034 0.0309
Deep Learning Methods (All-in-One Training)
PNN [13] 68.9 2.26 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
PanNet [18] 80.3 2.63 38.98 0.9468 0.0301 29.12 0.8927 0.0935 43.26 0.9668 0.0176
MSDCNN [19] 239.0 7.83 40.31 0.9580 0.0267 29.63 0.9033 0.0833 43.21 0.9671 0.0176
PanFlow [17] 87.3 2.86 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
PSCINN [15] 3321.5 108.84 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Pan-Mamba [8] 488.8 16.02 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
CFDCNet [11] 1700.8 55.73 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
@@ 3.2 Transformer Block
Figure 3: Quality comparison across SOTA methods on WV3 dataset. Refer to supplementary materials for more results.
Model QNR $\uparrow$
Separate All-in-One
MSDCNN [19] 0.7683 0.8898
PanNet [18] 0.7684 0.8726
PNN [13] 0.7527 0.8844
PSCINN [15] 0.7904 0.8849
PanTiny (Small) 0.7827 0.8751
PanTiny (Big) 0.7985 0.8793
PanFlow [17] 0.7910 0.8900
@@ 3.3 Loss Function
  • L1 Loss: We use the Charbonnier loss [4], a differentiable variant of the L1 norm that is less sensitive to outliers. For a batch of $B$ images with $N$ pixels each, it is defined as:
@@
  • SSIM Loss: To preserve perceptual quality and high-frequency structural details, we incorporate the Structural Similarity (SSIM) loss [16]. The SSIM index between two image patches $o$ and $g$ is:
@@
  • Focal-Inspired Regression Loss: Inspired by the principle of focal loss in classification, we adapt its core idea to regression to prioritize “hard” pixels. We term this $L_{Focal}$. Let $d_{i}=|O_{i}-G_{i}|$ be the absolute error for a given pixel $i$ (assumed to be in the [0, 1] normalized range). Our loss is formulated as:
@@
    where the constant $255$ scales the normalized error back to the 8-bit dynamic range before applying the exponent, stabilizing the gradient. In our experiments, we use a focusing parameter $r_{1}=1.1$. This formulation acts as a non-linear weighting function: by setting $r_{1}>1.0$, pixels with larger errors (hard” pixels) are polynomially up-weighted, compelling the model to focus on challenging details, similar in effect to a high-order $L_{p}$ loss.
@@
Through extensive experiments (detailed in Table 6 and our supplementary material), we determined the optimal weights to be $\lambda_{1}=1.5$, $\lambda_{2}=4.0$, and $\lambda_{3}=1.5$, which consistently delivered the best performance.
4 Experiments
Model Training WV2 WV3 GF2
[…]
Pan-Mamba [8] All-in-One 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
Separate 42.24 0.9729 0.0212 31.16 0.9299 0.0702 47.65 0.9894 0.0103
PNN [13] All-in-One 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
Separate 40.76 0.9624 0.0259 29.94 0.9121 0.0824 43.12 0.9704 0.0172
PanFlow [17] All-in-One 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
Separate 41.86 0.9712 0.0224 30.49 0.9221 0.0751 47.25 0.9884 0.0103
PSCINN [15] All-in-One 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Separate 41.85 0.9703 0.0223 30.56 0.9230 0.0748 47.11 0.9878 0.0107
CFDCNet [11] All-in-One 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
Separate 42.24 0.9733 0.0209 31.24 0.9327 0.0694 47.84 0.9902 0.0097
Pan-LUT [2] All-in-One - - - - - - - - -
Separate 39.84 0.9555 0.0286 28.82 0.8936 0.0935 42.66 0.9642 0.0189
Separate 42.16 0.9711 0.0217 30.61 0.9245 0.0747 48.93 0.9900 0.0092
4.1 Setup
Datasets. We conduct experiments on three public datasets: WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2). It is a standard community practice [13, 18] to pre-process the 8-band WV2 and WV3 data to 4-bands (R, G, B, NIR) to match GF2. We follow this standard, using the 4-band versions to ensure fair comparison and enable our “all-in-one” joint training. For our primary “all-in-one” experiments, we combine the training sets of all three. We follow standard protocols for evaluation, using both reduced-resolution and full-resolution test sets.
Evaluation Metrics. We provide a comprehensive evaluation using both reference and no-reference metrics. For reduced-resolution evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) [16], Spectral Angle Mapper (SAM) [20], and ERGAS [14]. For full-resolution evaluation, we use the no-reference metrics Dλ, Ds, and QNR [1].
Implementation Details. Our framework is implemented in PyTorch. All models are trained on a single NVIDIA RTX 4090 GPU. We use the ADAM optimizer with a learning rate of $5\times 10^{-4}$ and betas of $(0.9,0.999)$. A cosine annealing scheduler adjusts the learning rate over 500 epochs with a batch size of 16.
4.2 The “All-in-One” Training Paradigm
Our experimental evaluation is first designed to validate our core thesis: that the “all-in-one” training paradigm is a more robust and effective method for developing generalizable pan-sharpening models, solving the key challenges of FR generalization and deployment practicality.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
CFDCNet [11] 1700.8 55.73 42.50 0.9729 0.0205 31.11 0.9298 0.0707 49.07 0.9903 0.0091
Pan-Mamba [8] 488.8 16.02 41.77 0.9691 0.0226 30.36 0.9215 0.0769 45.84 0.9811 0.0134
PanFlow [17] 87.3 2.86 41.68 0.9688 0.0229 30.24 0.9197 0.0785 47.49 0.9865 0.0109
MSDCNN [19] 239.0 7.83 41.46 0.9669 0.0236 30.18 0.9189 0.0786 44.09 0.9730 0.0161
PNN [13] 68.9 2.26 40.84 0.9635 0.0256 29.82 0.9128 0.0834 43.40 0.9688 0.0173
PanNet [18] 80.3 2.63 40.79 0.9620 0.0256 29.82 0.9106 0.0841 43.76 0.9705 0.0167
4.2.1 Limitations of Traditional Paradigms
We first analyze the impact of our “all-in-one” paradigm on the traditional reduced-resolution (RR) benchmarks, as shown in Table 3. The results are mixed: compared to “Separate” models that are highly specialized (overfitted) to a single dataset, the “All-in-One” models (e.g., Pan-Mamba [8] and PSCINN [15]) show a significant performance drop on RR metrics. This highlights that the domain gaps between satellites are severe, and models must balance conflicting objectives. In contrast, our “PanTiny (Big)” model is remarkably robust, with only a minor drop on WV2/GF2 and almost no change on WV3. This demonstrates that simply chasing RR benchmark scores on a single dataset is not a robust strategy.
4.2.2 The True Value: Full-Resolution Generalization
The primary goal of pan-sharpening is real-world application, which requires generalizing from RR training to FR inference. Table 2 reveals the true, unambiguous value of our paradigm. For all models tested, switching from “Separate” to “All-in-One” training results in a dramatic and universal improvement in the full-resolution QNR metric. For instance, PanFlow’s [17] QNR score on WV2 jumps from 0.7910 to 0.8900, and MSDCNN’s [19] from 0.7683 to 0.8898. This is clear evidence that our paradigm trains models to be far more robust and generalizable to the real-world FR task. This, combined with the inherent benefit of deploying a single unified model (our Contribution 2), underscores the power of our proposed method.
4.3 PanTiny: An Efficient SOTA Model for the New Paradigm
Having established the “all-in-one” paradigm as a superior and more practical approach, we now benchmark models within this framework to validate our proposed PanTiny (Contribution 3).
4.3.1 Main Results on Multi-Dataset Training
Table 1 presents the main results. Our proposed PanTiny (Big) achieves SOTA performance across all three datasets, outperforming both classic and recent methods. Notably, it surpasses CFDCNet [11], a much larger model, on all metrics. PanTiny (Small) also achieves a highly competitive performance-to-efficiency balance. The results for PSCINN [15] highlight its instability in this multi-dataset setting, as it failed to complete training, which we consider a practical failure, further validating our robust design choice.
Model Params WV2 WV3 GF2
(K) PSNR SSIM PSNR SSIM PSNR SSIM
PanTiny (Small) 48.3 41.62 0.9685 30.38 0.9216 48.16 0.9884
PanTiny (Big) 81.7 41.85 0.9696 30.59 0.9238 48.61 0.9894
PanTiny (Large Body) 172.4 42.06 0.9708 30.67 0.9248 48.75 0.9896
PanTiny (Huge Body) 195.9 42.12 0.9711 30.74 0.9258 48.85 0.9898
4.3.2 Architectural Design and Efficiency
PanTiny’s SOTA performance is achieved with superior efficiency. As shown in Table 1, its 81.7K parameters and 2.68G FLOPs are orders of magnitude smaller than heavyweight competitors like CFDCNet (1700.8K params, 55.73G FLOPs) [11]. We further explore scaling in Table 5. By expanding our model to a ’Huge’ version (195.9K params), we can nearly match CFDCNet’s best-in-class performance on GF2 (48.85 vs 49.07 PSNR) at less than 12% of its parameter count. This reinforces our core argument that principled, efficient engineering surpasses brute-force scaling. Critically, PanTiny’s practical efficiency also lies in its reproducibility. Unlike Pan-Mamba [8] or CFDCNet [11] which may rely on specialized CUDA kernels or dependencies, PanTiny is built with standard PyTorch operators, ensuring it is easy to deploy and validate.